What is the focal length of a makeup mirror that produces a magnification of 1.90 when a person’s face is 12.0 cm away? (Make sure your focal length is positive if the mirror is concave, and negative if the mirror is convex.)
Answer #1

1/f = 1/do + 1/di

m = - di /do

1.9 = - di /12

di = -22.8 cm

1/f = 1/12 - 1/22.8

f = (22.8*12)/10.8

f = 25.33cm